Part A Find the frequency of light f radiated by an electron moving from orbit n1=2...

Part A

Find the frequency of light f radiated by an electron moving from orbit n1=2 to n2=1 inside of a He+ ion.

Express your answer in hertz to three significant figures.

Part B

In the Bohr model of hydrogen, the radius of the nth orbit is defined as

rn=a0n2Z,

where

a0=4???2mee2=5.29
Answer #1

En = -13.60 Z2/n2 eV
now Z = 2
En = -54.40 /n2 eV
a) E2 - E1 = 40.8 eV = hf
?f = 9.85*1015 Hz


b) r1 = a0/2 = 2.65*10-11 m